# Morphing Tokens Draw Strong Masked Image Models

### v2

@@ Abstract
Masked image modeling (MIM) is a promising option for training Vision Transformers among various self-supervised learning (SSL) methods. The essence of MIM lies in token-wise masked token predictions, with targets tokenized from images or generated by pre-trained models such as vision-language models. While tokenizers or pre-trained models are plausible MIM targets, they often offer spatially inconsistent targets even for neighboring tokens, complicating models to learn unified discriminative representations. Our pilot study confirms that addressing spatial inconsistencies has the potential to enhance representation quality. Motivated by the findings, we introduce a novel self-supervision signal called Dynamic Token Morphing (DTM), which dynamically aggregates contextually related tokens to yield contextualized targets. DTM is compatible with various SSL frameworks; we showcase an improved MIM by employing DTM, barely introducing extra training costs. Our experiments on ImageNet-1K and ADE20K demonstrate the superiority of our methods compared with state-of-the-art, complex MIM methods. Furthermore, the comparative evaluation of the iNaturalists and fine-grained visual classification datasets further validates the transferability of our method on various downstream tasks. Code is available at https://github.com/naver-ai/dtm.
@@ 1 Introduction
Since the success of Vision Transformer (ViT), numerous training strategies have arisen for ViT [15], including self-supervised learning (SSL) methods [10, 20, 18, 7]. Recent advances in masked image modeling (MIM) [52, 19, 4] have established it as a mainstream SSL approach for ViT. In essence, the crux of the MIM methods lies in leveraging token-wise optimization objectives derived from predicting masked tokens to given targets. Consequently, MIM methods have explored various approaches to generate effective target tokens, employing various pre-trained models including vision-language models [5, 34], utilizing momentum teachers [4, 56] or directly exploiting patchified images [19, 52].
@@
While tokenizers or pre-trained models have proven their potential as promising MIM targets [5, 34, 27, 49], we argue the pre-trained models tend to generate spatially noisy token representations of class information (e.g., inconsistent class labels), and their utilization as pre-training targets may incur negative impacts (See §3). For example, a pre-trained vision-language model reveals spatially inconsistent prediction results, as shown in Fig. 1. To delve deeper into the behaviors of token representations from pre-trained models, we analyze the impacts of spatially inconsistent representations from the perspective of discriminability and training capability. We initially observe reduced discriminability in zero-shot image classification and linear probing without token aggregation, also revealed with our continuous metric. Furthermore, another pilot distillation experiment reveals that the targets derived from a token aggregation method, not aware of context, disrupts pre-training. Unlikely, the model supervised by aggregated target tokens with context enjoys the potential of enhanced representation. Consequently, spatially inconsistent targets challenge learning one-to-one token maps, leading to suboptimal representation learning.
@@
Bearing those in mind, we introduce a novel token contextualization method Dynamic Token Morphing (DTM), where token morphing is a process that links contextually similar tokens and aggregates them for coherent and comprehensive representations. We conjecture that training should be accelerated by the guidance provided by composite representations of morphed tokens, derived from the aggregation of contextually related tokens (see Fig. 3). Specifically, we encode the token-wise target representations and derive matching relations among tokens through DTM. The merging process is performed for both tokens regarding the matching relation and aligning each morphed token with the corresponding morphed target tokens. The range of morphing can vary from a single token to the entire tokens, covering from token-wise to image-level representation learning. While there are many options, we simply opt for bipartite matching for morphing, achieving both efficiency and efficacy.
@@
Through extensive experiments, we verify our method’s general applicability and scalability. DTM improves fine-tuning accuracies as equipped with various SSL frameworks, affirming its general applicability as a basic element. Moreover, DTM consistently enhances the baselines from ViT-S/16 to ViT-L/16 with large margins, demonstrating its scalability. Comparison with previous SSL methods verifying its superiority. Our method showcases further transferability on the iNaturalist and fine-grained visual classification datasets.
@@ 2 Related Work
2.0.1 Masked image modeling.
Inspired by the promising performance of masked language modeling (MLM), BEiT [5] successfully extends MLM into the computer vision domain, using an external offline tokenizer from Dall-E [36]. iBOT [56] jointly trains the target encoder and the online tokenizer to remove the dependency on the external tokenizer. Data2vec [4] incorporates a momentum encoder to perform feature-level masked prediction tasks, leveraging representations from the multiple layers of neural networks. MAE [19] and SimMIM [52] demonstrate efficient masked image modeling by directly reconstructing masked input pixels without any tokenizer. On the other hand, several attempts have been made to exploit the pre-trained model as a tokenizer. BEiT v2 [34] pre-train a codebook for CLIP [35] to discretize a semantic space. MVP [48] exploits a tokenizer pre-trained with multimodal data to enhance the semantics for MIM. A line of studies [49, 38] using CLIP as a teacher to generate target representations have also been highlighted. Our method aims to utilize a teacher model more effectively, including CLIP, rather than just using it as a raw pre-trained model.
2.0.2 Token aggregation methods.
Token aggregation can conceptually be categorized as a token clustering method and usually aims for efficiency. Hard clustering methods like K-Means [30], K-Medoids [1], and Density-Peak Clustering with K-Nearest Neighbors (DPC-KNN) [23] enforce each data to belong to a single cluster exclusively. Bipartite matching [24] also aggregates data in a hard clustering manner, which optimizes pairs elements from two disjoint sets given objective function. Meanwhile, soft clustering is defined to let data belong to multiple clusters. LIT [33] employs deformable token merging layers to aggregate tokens between stages. Furthermore, some token pruning methods [37, 53, 40, 28] can be categorized into token aggregation methods; however, they intensely focused on compressing tokens to aim for a cost-efficient vision transformer. While the above token aggregation methods have mainly been employed to boost efficiency, our approach diverges significantly. We take the concept of token aggregation to address spatially noisy target tokens in token-level supervision, thereby enhancing the efficacy of MIM in terms of precision.
@@ 3 Pilot Study
In this section, we study the behavior of token representations generated by pre-trained models and their negative impacts on pre-training.
Spatial Inconsistency.
We begin by visualizing token representations from a vision-language model [35] to illustrate spatial inconsistency.
The inconsistent class predictions observed on each patch, as depicted in Fig. 1, reveal noisy token representations.
Impact of Spatial Inconsistency.
The spatially noisy token representations has the potential to lead to an underestimation of the capabilities of the vision-language models.
To understand the extent of these effects, we quantitatively assess various metrics at a dataset scale in Table 1. The quantitative analyses reveal that enhancing spatial coherence in the token representations results in improved performance, thus highlighting the importance of addressing spatial inconsistency.
Spatial Inconsistency in Representation Learning. The lack of spatial coherence in token representations have substantial negative impacts on the representation learning. Utilizing such noisy representations as targets can undermine the capabilities of the models. We employ representation learning to assess token aggregation methods to confirm the efficacy of handling spatial inconsistency via a refined context-aware method.
@@ 3.1 Spatial Inconsistency
We define the spatial inconsistency among token representations with an illustrative example. Given the input image in Fig. 1(a), we visualize the token-wise predictions without and with token aggregation in Fig. 1(b) and 1(c), respectively. Correct and incorrect tokens are marked in green and red, respectively, with a gradient to darker shades of red, indicating a more significant deviation from the true class. Despite the proximity and contextual similarity among tokens, wrong tokens in the green box (bottom right) in Fig. 1(b) exhibit spatially inconsistent prediction results (French Bulldog, Gorilla, Bull Mastiff, Hotdog, and Newfoundland Dog) while tokens in Fig. 1(c) show correct or relatively consistent predictions (french bulldog). Moreover, token aggregation for predictions yields improved accuracy, with 113 correct tokens out of 196 and a gap of 31 tokens compared to the counterpart. The comparison between the two prediction results highlights the spatial inconsistency among tokens from a pre-trained model, which could potentially disrupt representation learning when used as targets.
3.2 Impact of Spatial Inconsistency
@@
To quantitatively assess the impact of spatial inconsistency at a dataset scale, we compute the token-wise ensemble prediction using a global pool, with and without token aggregations. Fundamentally, we predict class scores for each token and average these scores across all tokens within the given image. When predicting with token aggregation, we group semantically relevant tokens and average their representations group-wise prior to token-wise ensemble prediction. We employ CLIP-B/16 [35] for our study, so we aggregate 98 tokens for token aggregation, which occupy half of the entire tokens.
@@
Table 1 reports the results of zero-shot image classification, linear probing, and averaged cosine similarity between the [CLS] on ImageNet-1K. Note that we do not perform any extra training here. Our results show that using aggregated tokens consistently exceeds the performances of those without token aggregation across all metrics. This suggests that aggregating contextually related tokens can address inconsistencies and lead to performance gains.
@@
Note that the final metric we use is an alternative: a patch-wise continuous similarity metric. We measure the cosine similarity between the [CLS] token and each patch, which is also a patch-wise metric but continuous111 We further involve a continuous metric to assess spatial inconsistency, as it tracks network responses continuously compared with discrete classification metrics.. As observed in Table 1, the averaged similarities are computed to 0.56 vs. 0.53 for each case. This new metric, directly linked to the [CLS] token, suggests a more direct relationship between improved accuracy and reduced spatial inconsistency. The trend aligns with the other discrete metric, depicting patch-wise classification of 30.8 and 26.5 with and without token aggregation, respectively. The quantitative assessments, using both continuous and discontinuous metrics, demonstrate that token representations exhibit spatial inconsistency. Taking this further, we believe more effectively addressing inconsistency would yield a greater impact.
@@ 3.3 Spatial Inconsistency in Representation Learning
We now study the impact of token aggregation on representation learning. This analysis aims to highlight the spatial inconsistency in self-distillation signal could potentially diminish the quality of representations. Our comparison ranges distillations with the signals of 1) token-wise, 2) downsampling, 3) layer-wise bipartite matching during encoding, 4) superpixel, and 5) token morphing via bipartite matching. Note that token morphing will be introduced in §4.1, but we show its effectiveness in advance. We pre-train ViT-B/16 for 50 epochs and perform linear probing for 50 epochs on ImageNet-1K for each configuration.
First, Table 2 exhibits distillation with token aggregation methods such as token morphing via bipartite matching, and superpixel clustering outperforms that of the token-wise distillation approach. Once again, this reveals that token representations indeed exhibit spatial inconsistency, as we claimed. Furthermore, token aggregation methods for contextually related tokens could manage this inconsistency. Since the downsampling approach aggregates tokens without context awareness, and the layer-wise bipartite matching approach risks harming intermediate representations during encoding, both approaches exhibit diminished returns. The findings highlight the significance of employing context-aware token aggregations that preserve token count while addressing spatial inconsistency.
@@
4 Method - Dynamic Token Morphing
We observed that 1) pre-trained models often produce noisy, spatial inconsistency token-level targets for learning, closely linked with performance degradation; 2) token aggregation methods could handle spatial inconsistency but are insufficient as a supervisory signal; 3) a well-designed method is needed to consider context and reduce noise more effectively. Motivated by the observations, we introduce an advanced token aggregation method Dynamic Token Morphing (DTM) for masked image modeling and self-supervised pre-training. In DTM, we contextually aggregate tokens to derive random numbers of morphed tokens aiming to encapsulate diversified semantic information. The core idea of DTM is illustrated in Fig. 3, where the DTM module is straightforwardly added to the baseline MIM scheme. DTM is demonstrated for learning conventional token-level MIMs by aligning morphed tokens from online and target encoders.
@@ 4.1 Preliminary
To this end, the input of the encoders becomes where for and for .
@@
Token morphing. Differ to token aggregation methods, which prioritize efficiency via token reduction, token morphing specifically targets the spatial inconsistency in token representations by taking context into account. It connects contextually relevant tokens to smooth them without reducing the number of tokens (i.e., maintaining the token count). We define the process with the token morphing function based on the morphing schedule , a sequence of token numbers to morph. Here, is a general notation for a function that calculates similarity using a matching algorithm (e.g., bipartite matching or K-Means clustering) and returns a token morphing matrix , where means the number of token groups after morphing. Eventually, we can morph tokens following Eq. 3.
4.2 Overall Pipeline of Dynamic Token Morphing

### v3

@@ Abstract
Masked image modeling (MIM) has emerged as a promising approach for training Vision Transformers (ViTs). The essence of MIM lies in the token-wise prediction of masked tokens, which aims to predict targets tokenized from images or generated by pre-trained models like vision-language models. While using tokenizers or pre-trained models are plausible MIM targets, they often offer spatially inconsistent targets even for neighboring tokens, complicating models to learn unified and discriminative representations. Our pilot study identifies spatial inconsistencies and suggests that resolving them can accelerate representation learning. Building upon this insight, we introduce a novel self-supervision signal called Dynamic Token Morphing (DTM), which dynamically aggregates contextually related tokens to yield contextualized targets, thereby mitigating spatial inconsistency. DTM is compatible with various SSL frameworks; we showcase improved MIM results by employing DTM, barely introducing extra training costs. Our method facilitates training by using consistent targets, resulting in 1) faster training and 2) reduced losses. Experiments on ImageNet-1K and ADE20K demonstrate the superiority of our method compared with state-of-the-art, complex MIM methods. Furthermore, the comparative evaluation of the iNaturalists and fine-grained visual classification datasets further validates the transferability of our method on various downstream tasks. Code is available at https://github.com/naver-ai/dtm.
@@ 1 Introduction
Since the success of Vision Transformers (ViTs) (Dosovitskiy et al., 2021), numerous training strategies have been developed for ViTs, including self-supervised learning (SSL) methods (Chen et al., 2020; He et al., 2019; Grill et al., 2020; Caron et al., 2021). Recent advances in masked image modeling (MIM) (Zhou et al., 2022; He et al., 2022; Peng et al., 2022; Baevski et al., 2022; Heo et al., 2023; Kim et al., 2024) have solidified its position as a primary SSL approach for ViT. The crux of the MIM methods is leveraging token-wise optimization objectives by predicting masked tokens to match given targets. MIM methods have explored various approaches to assign effective target tokens, employing various pre-trained models including vision-language models (Bao et al., 2021; Peng et al., 2022), utilizing momentum encoders (Baevski et al., 2022; Zhou et al., 2022) or directly exploiting patchified images (He et al., 2022; Xie et al., 2022).
@@
While tokenizers or pre-trained models have proven effective as MIM targets (Bao et al., 2021; Peng et al., 2022; Li et al., 2022b; Wei et al., 2022c), we argue they often generate spatially noisy token representations of class information (e.g., inconsistent class labels), which may incur negative impacts when utilized as pre-training targets; for example, a pre-trained vision-language model exhibits spatially inconsistent prediction results for tokens, as shown in Fig. 1. To further explore the behavior of token representations from pre-trained models, we analyze the effects of spatially inconsistent representations from the perspective of the models’ capability. Our pilot exploration shows reduced discriminability in zero-shot classification and continuous similarity metrics without token aggregation. Another pilot distillation experiment reveals that targets by context-unaware token aggregation disrupt pre-training. As a result, spatially inconsistent targets challenge learning one-to-one token maps, leading to suboptimal representation learning.
@@
Bearing this in mind, we introduce a novel token contextualization method called Dynamic Token Morphing (DTM), where token morphing links contextually similar tokens and aggregates them to produce coherent and comprehensive representations. We conjecture that training can be accelerated through the guidance of composite representations of morphed tokens derived from the aggregation of contextually related tokens. Specifically, we encode the token-wise target representations and derive matching relations among tokens using DTM. The merging process is applied to both online and target tokens regarding the matching relation and aligns each morphed token with the corresponding morphed target token. The range of morphing can vary from a single token to all tokens, covering from token-wise to image-level representation learning. Among various options, we opt for bipartite matching for morphing, achieving both efficiency and efficacy.
@@
Through extensive experiments, we verify our method’s general applicability and scalability. DTM could improve fine-tuning accuracies on ImageNet (Russakovsky et al., 2015) classification and ADE20K (Zhou et al., 2017) segmentation when equipped with various SSL frameworks for pre-training on ImageNet, demonstrating its versatility as a fundamental component. Comparison with previous SSL methods verifies its superiority. Our method showcases further transferability on the iNaturalist (Van Horn et al., 2018) and fine-grained visual classification datasets (Horn et al., 2015; Krizhevsky, 2009; Khosla et al., 2011).
@@ 2 Related Work
Masked image modeling. Inspired by the promising performance of masked language modeling (MLM), BEiT (Bao et al., 2021) successfully extends MLM into the computer vision domain, using an external offline tokenizer from Dall-E (Ramesh et al., 2021). iBOT (Zhou et al., 2022) jointly trains the target encoder and the online tokenizer to remove the dependency on the external tokenizer. Data2vec (Baevski et al., 2022) incorporates a momentum encoder to perform feature-level masked prediction tasks, leveraging representations from the multiple layers of neural networks. MAE (He et al., 2022) and SimMIM (Xie et al., 2022) demonstrate efficient masked image modeling by directly reconstructing masked input pixels without any tokenizer. On the other hand, several attempts have been made to exploit the pre-trained model as a tokenizer. BEiT v2 (Peng et al., 2022) pre-trains a codebook for CLIP (Radford et al., 2021) to discretize a semantic space. MVP (Wei et al., 2022b) exploits a tokenizer pre-trained with multimodal data to enhance the semantics for MIM. A line of studies (Wei et al., 2022c; Ren et al., 2023) using CLIP as a teacher to generate target representations have also been highlighted. Our method aims to utilize a teacher model more effectively, including CLIP, rather than just using it as a raw pre-trained model.
Token aggregation methods. Token aggregation can conceptually be categorized as a token clustering method and usually aims for efficiency. Hard clustering methods like K-Means (Lloyd, 1982), K-Medoids (kme, 1990), and Density-Peak Clustering with K-Nearest Neighbors (DPC-KNN) (Jiang et al., 2019) enforce each data to belong to a single cluster exclusively. Bipartite matching (Karp et al., 1990) also aggregates data in a hard clustering manner, which optimizes pairs elements from two disjoint sets given objective function. Meanwhile, soft clustering is defined to let data belong to multiple clusters. LIT (Pan et al., 2022) employs deformable token merging layers to aggregate tokens between stages. Furthermore, some token pruning methods (Rao et al., 2021; Xu et al., 2022; Tang et al., 2022; Liang et al., 2022) can be categorized into token aggregation methods; however, they intensely focused on compressing tokens to aim for a cost-efficient vision transformer. While the above token aggregation methods have mainly been employed to boost efficiency, our approach diverges significantly. We take the concept of token aggregation to address spatially noisy target tokens in token-level supervision, thereby enhancing the efficacy of MIM in terms of precision.
@@ 3 Pilot Study
In this section, we study the behavior of token representations generated by pre-trained models and their negative impacts on pre-training. Our study can be summarized as follows:
Revealing spatial inconsistency and its impact (§3.1). We begin by visualizing token representations from a pre-trained vision-language model to define spatial inconsistency: as shown in Fig. 1, the inconsistent class predictions across different patches suggest spatially noisy token representations. We further investigate impact of spatial inconsistency, as relying on the noisy token representations may underutilize the pre-trained model’s capability. The quantitative analyses in Table 1 show that enhancing spatial coherence in the token representations results in improved performance, which highlights the importance of addressing spatial inconsistency.
Spatial inconsistency in representation learning (§3.2). Our study shifts to a practical focus on pre-training for representation learning under supervision (i.e., target). We first reveal that the lack of spatial coherence in token representations significantly hinders learning and using such noisy representations as targets can weaken the supervision model. We then assess token aggregation methods as the supervision to confirm the noticeable efficacy of a context-based method for handling spatial inconsistency.
@@ 3.1 Spatial Inconsistency
Given the input image in Fig. 1(a), we visualize the token-wise zero-shot classification results without and with token aggregation in Fig. 1(b) and 1(c), respectively. Correct and incorrect tokens are marked in green and red, respectively, with a gradient to darker shades of red, indicating a more significant deviation from the true class. Despite the proximity and contextual similarity among tokens, wrong tokens in the green box (bottom right) in Fig. 1(b) exhibit spatially inconsistent prediction results (French Bulldog, Gorilla, Bull Mastiff, Hotdog, and Newfoundland Dog) while tokens in Fig. 1(c) show correct or relatively consistent predictions (French Bulldog). Moreover, token aggregation for predictions yields improved accuracy, with 113 correct tokens out of 196 and a gap of 31 tokens compared to the counterpart. This highlights the spatial inconsistency among tokens from a pre-trained model, which could potentially disrupt representation learning when used as supervision.
@@
Impact of spatial inconsistency. To quantitatively assess the impact of spatial inconsistency at a dataset scale, we compute the token-wise ensemble prediction using a global pool, with and without token aggregations. Fundamentally, we predict class scores for each token and average these scores across all tokens within the given image. When predicting with token aggregation, we group semantically relevant tokens and average their representations group-wise prior to token-wise ensemble prediction. We employ CLIP-B/16 (Radford et al., 2021) for our study, so we aggregate 98 tokens for token aggregation, which occupy half of the entire tokens.
@@
Table 1 below reports the results of zero-shot image classification, linear probing, and averaged cosine similarity with the [CLS] token on ImageNet-1K. Note that we do not perform any extra training here. Our results show that using aggregated tokens consistently exceeds the performances of those without token aggregation across all metrics. This suggests that aggregating contextually related tokens can address inconsistencies and lead to performance gains.
@@
Note that the final metric we use is a continuous metric that computes patch-wise similarity. Specifically, it measures the cosine similarity between the [CLS] token and each patch, which is also a patch-wise metric but continuous111 We believe a continuous metric to assess spatial inconsistency could track network responses continuously compared with the discrete classification metrics.. As observed in Table 1, the averaged similarities are computed to 0.56 vs. 0.53 for each case. This new metric, directly linked to the [CLS] token, suggests a more direct relationship between improved accuracy and reduced spatial inconsistency. The trend aligns with the other discrete metric, depicting patch-wise classification of 30.8 and 26.5 with and without token aggregation, respectively. The quantitative assessments, using both continuous and discontinuous metrics, demonstrate that token representations exhibit spatial inconsistency. Taking this further, we believe more effectively addressing inconsistency would yield a greater impact.
@@ 3.2 Spatial Inconsistency in Representation Learning
| Distillation signals | Linear prob () | Speed↓ (ms/img) |
| (1) Token-wise (baseline) | 70.9 | 0.629 |
| (2) Downsampling | 68.8 (-2.1) | 0.647 |
| (3) Layer-wise Bipartite Matching (Bolya et al., 2022) | 70.2 (-0.7) | 0.633 |
| (4) Superpixel clustering (Achanta et al., 2012) | 72.6 (+1.7) | 0.740 |
| (5) Token morphing via Bipartite Matching | 72.2 (+1.3) | 0.653 |
We now study token aggregations for handling spatial inconsistency in representation learning. This analysis aims to highlight the spatial inconsistency in supervision signals (i.e., self-distillation or pre-training targets) that potentially diminish the quality of representations. Our comparison ranges distillations with the signals of 1) token-wise, 2) downsampling, 3) layer-wise bipartite matching during encoding, 4) superpixel, and 5) token morphing via bipartite matching. Note that token morphing will be introduced in §4.1, but we show its effectiveness in advance. We pre-train ViT-B/16 for 50 epochs and perform linear probing for 50 epochs on ImageNet-1K for each configuration.
Table 2 exhibits distillation with context-aware token aggregation methods such as superpixel clustering and bipartite matching-based token morphing outperforms that of the token-wise distillation approach. We first observe token representations indeed exhibit spatial inconsistency, as we claimed. Furthermore, token aggregation methods for contextually related tokens could manage this inconsistency. Since the downsampling approach aggregates tokens without context awareness, and the layer-wise bipartite matching approach risks harming intermediate representations during encoding, both approaches exhibit diminished returns. The findings highlight the significance of employing context-aware token aggregations that preserve token count while addressing spatial inconsistency.
@@
4 Method
We observed that 1) supervision signal by a pre-trained model often produces noisy and spatially inconsistent token-level targets for learning, closely linked with performance degradation; 2) naive token aggregation methods could partially handle spatial inconsistency but are insufficient as a supervisory signal; 3) a well-designed method is favorable for considering context and reducing noise more effectively. Motivated by the observations, we introduce an advanced token aggregation method called Dynamic Token Morphing (DTM) for masked image modeling and self-supervised learning. DTM contextually aggregates tokens to derive random numbers of morphed tokens aiming to encapsulate diversified semantic information. The core idea of DTM is illustrated in Fig. 3, where the DTM module is straightforwardly added to the baseline (e.g., MIM). DTM encourages conventional token-level MIMs by aligning morphed tokens from online and target encoders by reducing spatial inconsistency concerning context.
@@ 4.1 Preliminary
To this end, the input of the encoders becomes , where for and for .
@@
Token morphing. Differing from token aggregation methods, which prioritize efficiency via token reduction, token morphing specifically targets the spatial inconsistency in token representations by taking context into account. It connects contextually relevant tokens to smooth them without reducing the number of tokens (i.e., maintaining the token count). We define the process with the token morphing function based on the morphing schedule , a sequence of token numbers to morph. Here, is a general notation for a function that calculates similarity using a matching algorithm (e.g., bipartite matching or K-Means clustering) and returns a token morphing matrix , where means the number of token groups after morphing and denotes the feature dimension. Eventually, we can morph tokens following Eq. 3.
4.2 Overall Dynamic Token Morphing Pipeline